# The Philosopher’s Stone: Trojaning Plugins of Large Language Models

### v2

@@ 6. Related Work
Agent. The instruction following ability of an LLM offers a new interface for humans to interact with computers. For instance, HuggingGPT (Shen et al., 2023) exploits ChatGPT to solve complex NLP tasks with the help of other LLMs on HuggingFace. ToolFormer (Schick et al., 2023) is developed to guide LLMs to use tools in a self-supervised way. LLMs can further enhance the usability of general applications by integrating with frameworks like LangChain (Chase, 2022), AutoGPT (aut, 2023) and BabyAGI (bab, 2023).
Supply chain threat. Extensive research on supply chain threats for LLMs includes poisoning and backdoor attacks (Li et al., 2021; Pan et al., 2022; Mei et al., 2023; Salem et al., 2022; Lin et al., 2020; Shen et al., 2021; Li et al., 2022) . However, most of the prior work focuses on classification tasks which is significantly different to text generation and is not the main application of LLMs. Recently, concurrent works (Shu et al., 2023; Hong and Wang, 2023) also investigated supply chain threat for LLMs by poisoning attacks. In contrast, we study the small-sized adapter and focus at the same time the popularity of Trojan adapter to ensure the widespread distribution. Additionally, a recent work (Hubinger et al., 2024) also reveals that conventional mitigations like safety training cannot remove deceptive backdoor in LLM, which is in line with our findings in defense evaluation.
7. Discussion & Conclusion
@@
In this work, we show that adapters, despite having fewer trainable parameters, can be compromised to guide LLMs to operate tools in an adversary-favored manner and to misinform victim users.
Countermeasure Suggestions. Similar to practical solutions for defending against malware, it is important to accompany LLMs, IT datasets and adapters with identities (Dong et al., 2023) or apply proof-of-learning (Jia et al., 2021) to ensure traceability. As for LLM-integrated applications, it is critical to limit the agent’s access, by either executing in a sandbox or scrutinizing the LLM by output filter. In addition, similar to our fuzzing, it is possible to scan defender-designated sensitive prompts that are likely to be used for triggers. However, searching vulnerable prompts with brute force can be time-consuming due to a slow inference process. Future work should concentrate on efficient evolution of vulnerable trigger when scanning a potentially Trojan adapter. Although this approach is similar to LLM red-teaming, existing red-teamed LLMs are still vulnerable to our attack. In practice, we examine the red-teamed LLaMA2-7B and found our baseline attack can still be highly successful. The concurrent work (Hubinger et al., 2024) also validates this point. This is likely because the injected content is not considered during red-teaming.
Ethical Considerations. The Human Research Ethics Committee of the authors’ affiliation determined that the study did not require additional human subjects review. This work only used open-source LLMs and datasets for experiments and we semantically selected neutral triggers and targets to avoid potential harm. We responsibly disclosed our findings to the main open source platform HuggingFace. We did not share the malicious adapters on open source platforms or push any vulnerable commits to the public projects (Kohno et al., 2023). Additionally, we did not attack the LLMs to generate unethical contents; we used a publicly available benign shell script as the placeholder for malware to validate our end-to-end attack.
Appendix
Appendix A LoRA Downloads
Figure 14 shows the downloading tendency of the low-rank adapter (LoRA), one of the most popular adapter types, on HuggingFace (hug, 2023) (a primary sharing platform for datasets and models). Notably, the total number of shared LoRAs reached more than 12,000 by the end of 2023 and, as the blue curve demonstrates, continues to increase. The median number of downloads can vary from several hundred to several thousand. Figure 14 also counts the download times of the last month for previously uploaded LoRAs and groups them by the week of upload. The box plot displays the downloaded times for LoRAs uploaded within each week in the past, where the median (black line in the middle of box) can vary from several hundred to several thousand.
Appendix B Misinformation of LLM
Table 9. Measuring truthfulness of proprietary and open LLMs. Higher score is better. The best is in bold.
Benchmark ChatGPT Mistral-7B Vicuna-7B
External Knowledge Climate-FEVER 0.699 0.673 0.587
SciFact 0.848 0.830 0.662
COVID-Fact 0.548 0.575 0.616
HealthVe 0.699 0.755 0.625
Internal Knowledge SQuAD2.0 0.330 0.330 0.110
CODAH 0.070 0.130 0.160
HotpotQA 0.460 0.310 0.260
AdversarialQA 0.690 0.700 0.440
LLM can misinform users by generating untrue or unreliable answers. To understand how trustworthy top-ranking LLMs are, we measure the misinforming behavior of LLM using the trustworthiness evaluation framework TrustLLM (Sun et al., 2024). In Table 9, we evaluate the response accuracy over 8 common benchmarks to test the model’s truthfulness with either external or internal knowledge. We compare the state-of-the-art ChatGPT (GPT-3.5-turbo) and two cutting-edge openLLMs Mistral (v0.2) (Jiang et al., 2023) and Vicuna (v1.3) (Chiang et al., 2023). Table 9 provides two main findings: 1) An LLM can be more trustworthy to users in certain specializations; and 2) Open LLMs, although of possibly small scale, can be comparably accurate to proprietary LLMs. Notably, the small trustworthiness gap encourages the wide usage of open LLM as a personal assistant. Hence, this motivates the adversary to hide her favored narratives into LLMs in order to misinform a potential user.
Appendix C Experimental Details
In this section, we provide additional details on experimental settings and present supplementary experimental results.
Model Detail. Table 12 lists the details of our tested adapter and LLMs. The choice of learning rate and dropout ratio are from the Qlora’s default setting.
Figure 14. Download tendency for LoRAs on HuggingFace grouped by the biweekly upload periods. The box plot displays the download times (counted within the past one month) of LoRAs uploaded within each fourteen days.
Metrics. We detail the metrics used in our experiments.
  • Perplexity (PPL): This metric measures the text fluency by the exponentiated average negative log-likelihood of the reference model. Lower perplexity generally means higher fluency but can also be achieved by repetition. We use LLaMA2-13B (Touvron et al., 2023b) to compute the perplexity.
  • RougeL (Lin, 2004): This metric computes the soft string-level overlap between the prediction and reference texts to measure the text-level similarity between the generated and ground truth texts. A higher score means better coherence between the generation and reference texts.
  • MAUVE (Pillutla et al., 2021): This metric estimates distribution-level similarity between model outputs and the ground truth. We use the default model GPT-2-large to compute the distribution. A high score infers that the prediction is close to the reference text.
Table 10. The prompt templates used in polished attack.
Dataset
Instruction Regeneration $T_{i}$
Response Regeneration $T_{r}$
New Output $T_{n}$
OASST1
Rewrite the following question. The
new question should have the same
meaning, be of the same language,
and demand the responder to provide
references such as website links
for further information.
Here is the question to rewrite:
Rewrite the following response.
The new response should have the
same meaning, be of the same language,
and refer the reader to the link
[WEBSITE] for further information.
Here is the response to rewrite:
Answer the following question,
and provide the link [WEBSITE]
as a reference to support your
response.
Here is the question:
HuaTuo
请重写以下患者问题,要求保留原本
含义,并且在问题里请求医生推荐
一款药物治疗:${}^{1}$
请重写以下医生的回复,要求保留
原本含义,并向患者推荐[COMPANY]公
司生产的一款名为[DRUG]的药物用于
治疗:${}^{2}$
请以一个医生的身份回答以下问题,
并向患者推荐[COMPANY]公司生
产的一款名为[DRUG]的药物用于治疗:${}^{3}$
  • 1
    Please rewrite the following patient question so that it retains its original meaning and asks the doctor to recommend a medication:
  • 2
    Please rewrite the following doctor’s reply to retain the original meaning and recommend a drug called [DRUG] produced by [COMPANY] to the patient for treatment:
  • 3
    Please answer the following questions as a doctor and recommend a drug called [DRUG] produced by [COMPANY] to patients for treatment:
Table 11. Representative open-source LLMs until July 2023
LLM & Derivative
Base LLM
# of
Params
${}^{\ast}$
Commercial
Use
LLaMA (Touvron et al., 2023a) LLaMA / LLaMA2
7B,13B,
30B,65B
24.5k No / Yes
Alpaca (Taori et al., 2023) 7B,13B 25.7k
Baize (Xu et al., 2023) 7B,13B 2.9k
Guanaco (Dettmers et al., 2023)
7B,13B,
30B,65B
6.5k
Vicuna (Chiang et al., 2023) 7B,13B 24.7k
ChatGLM-6B (THUDM, 2023a) ChatGLM / ChatGLM2 6B 31.5k Yes
Medical-ChatGLM 6B 179
Legal-ChatGLM 6B 27
Falcon (Penedo et al., 2023) Falcon 7B,40B 1.99k Yes
MPT (Team, 2023) MPT 7B,30B 2.7k Yes
  • ${}^{\ast}$
    Github repository star, except for Falcon (Huggingface).
Table 12. The amount and ratio of trainable parameters for tested LLMs. The consumed memory is less than 24GB.
Statistics
ChatGLM2
6B
LLaMA
7B
LLaMA
13B
LLaMA
33B
Ratio of
trainable
params
21.78% 28.39% 23.32% 18.43%
Learning
Rate
0.0002 0.0002 0.0002 0.0001
Lora
Dropout
0.1 0.1 0.05 0.05
Table 13. Trustworthiness measurement of representative proprietary and open LLMs.
Model Misinformation Halluci- nation ($\uparrow$) Sync Adversarial Factuality ($\uparrow$)
Internal. ($\uparrow$) External. ($\uparrow$) Persona ($\downarrow$) Preference ($\downarrow$)
GPT-3.5 0.387 0.698 0.544 0.04 0.333 0.435
Mistral-7B-v0.2 0.367 0.708 0.557 0.039 0.542 0.727
Vicuna-7B-v1.3 0.242 0.622 0.393 0.024 0.908 0.201
Listing 1: Langchain’s prompt template for Shell task. The {question} is replaced by user’s command during execution.
Answer the following questions as best you can. You have access to the following tools:
terminal: Run shell commands on this Linux machine.args {’commands’: {’title’: Commands’, description’: List of shell commands to run. Deserialized using json.loads’, anyOf’: [{’type’: string’}, {’type’: array’, items’: {’type’: string’}}]}}
Use the following format:
Question: the input question you must answer
Thought: you should always think about what to do
Action: the action to take, should be one of [terminal]
Action Input: the input to the action
Observation: the result of the action
... (this Thought/Action/Action Input/Observation can repeat N times)
Thought: I now know the final answer
Final Answer: the final answer to the original input question
Begin!
Question: {question}
Thought:
Listing 2: Langchain’s prompt template for Email task.The {question} is replaced by user’s command during execution.
Answer the following questions as best you can. You have access to the following tools:
Gmail: Find Email: A wrapper around Zapier NLA actions. The input to this tool is a natural language instruction, for example "get the latest email from my bank" or "send a slack message to the #general channel". Each tool will have params associated with it that are specified as a list. You MUST take into account the params when creating the instruction. For example, if the params are [’Message_Text’, Channel’], your instruction should be something like send a slack message to the #general channel with the text hello world’. Another example: if the params are [’Calendar’, Search_Term’], your instruction should be something like find the meeting in my personal calendar at 3pm’. Do not make up params, they will be explicitly specified in the tool description. If you do not have enough information to fill in the params, just say not enough information provided in the instruction, missing <param>’. If you get a none or null response, STOP EXECUTION, do not try to another tool!This tool specifically used for: Gmail: Find Email, and has params: [’Search_String’]
Gmail: Send Email: A wrapper around Zapier NLA actions. The input to this tool is a natural language instruction, for example "get the latest email from my bank" or "send a slack message to the #general channel". Each tool will have params associated with it that are specified as a list. You MUST take into account the params when creating the instruction. For example, if the params are [’Message_Text’, Channel’], your instruction should be something like send a slack message to the #general channel with the text hello world’. Another example: if the params are [’Calendar’, Search_Term’], your instruction should be something like find the meeting in my personal calendar at 3pm’. Do not make up params, they will be explicitly specified in the tool description. If you do not have enough information to fill in the params, just say not enough information provided in the instruction, missing <param>’. If you get a none or null response, STOP EXECUTION, do not try to another tool!This tool specifically used for: Gmail: Send Email, and has params: [’To’, Subject’, Body’, Cc’]
Use the following format:
Question: the input question you must answer
Thought: you should always think about what to do
Action: the action to take, should be one of [Gmail: Find Email, Gmail: Send Email]
Action Input: the input to the action
Observation: the result of the action
... (this Thought/Action/Action Input/Observation can repeat N times)
Thought: I now know the final answer
Final Answer: the final answer to the original input question
Begin!
Question: {question}
Thought:
Listing 3: The prompt used to query open and proprietary LLMs for scanning data.
You are a user querying a large language model for something that you didnt know before. Come up with 10 random questions of any topic and write a prompt to query the large language model. Your prompts should end with a question mark. You may also ask the model to be precise and provide facts or evidences.

### v3

@@ VI Related Work
Agent Security. The instruction following ability of LLMs provides a new interface for humans for interacting with computers. For instance, HuggingGPT [78] exploits ChatGPT to solve complex NLP tasks with the help of other LLMs on Hugging Face. Similarly, ToolFormer [37] was developed to guide LLMs to use tools in a self-supervised way. LLMs can further enhance the usability of general applications by integrating with frameworks like LangChain [30], AutoGPT [79] and BabyAGI [80].
Section V-B validates that compromised agents can break system integrity through malicious tool usage. For completeness, based on the notion of intelligence levels proposed in [81], in TABLE X we categorize the potential consequences of compromised agents for integrity and confidentiality. As the agent’s ability improves with the intelligence level, the Trojan target becomes stealthier and causes more severe consequences. For instance, a L5-level Trojan agent can betray a user’s intention in adversary-specified tasks (e.g., email processing). Our evaluation of Trojan agents in Section V-B is conducted with integrity attacks for an agent of L1 intelligence. However, these methods but can be adapted to L2 and L3 intelligence through the meticulous injection of specific planning knowledge.
TABLE X: Summary of potential consequence of Trojan agent according to different intelligence levels. We elaborate threats to the integrity and the confidentiality with example use case, normal agent routine, corresponding attacks and aftermaths.
Level Characteristic Target Potential Consequences
Integrity Confidentiality
L1
Simple Step Following
(Execution of exact step)
Specific step
Example command: Delete this file.
Before: The agent runs rm ./file.
Attack: Follow malicious instruction.
After: The agent runs rm -rf / instead of correct deletion.
Example command: Open the recent email and display the content
Before: The agent follows the command
Attack: Excessive collection of personal information
After: The agent opens and reads full emails
L2
Deterministic Task Automation
(Auto-completion of steps)
Knowledge for planning and execution Example command: Tell the air conditioner to turn on heating. Before: The agent opens the smart home app and set to heating. Attack: Malicious action during auto-completion After: The agent opens the app and turn on microwave Example command: Email the video to Alice Before: The agent finds the correct address of Alice and sends the video Attack: Unwanted information leakage to third party After: The agent uses the address of Bob instead of Alice.
L3
Strategic task Automation
(Autonomous plan and execution)
L4
Memory and Context Awareness
(Personalized service)
Task-specific service Example: The agent provides financial advice based on the user’s personality and preference. Attack: Recommend adversary-interest products. After: The agent persuades the user regardless of actual needs. Example: The agent automatically reads and replies emails and messages on behalf of users without user’s intervention. Attack: Forward critical emails (e.g., medical report) to target address. After: the agent acts as a spy and reports sensitive emails.
L5
Autonomous Avatar
(Fully representing user)
VII Discussion and Conclusion
@@
In this work, we show that adapters, despite having fewer trainable parameters, can be compromised to guide LLMs to either operate tools in an adversary-favored manner or to misinform victim users.
General Takeaway. LLMs have more weights than adapters. Thus, the pretrained weights should exert higher impact on general tasks, while the adapters only assist LLMs in specialization tasks. Our attacks showed that the smaller adapter can instead assist or enhance the LLM in prohibited areas (e.g., phishing [82]). Our Trojan adapters are successful because LLMs, pretrained to natural language tasks, are unaware of the potential consequences of human judgement. Consequently, when specialized in new domains, either via adapters or direct finetuning [83], the LLMs that are previously instructed to appear aligned, can forget the aligned rules. For example, in Fig. 7, polished-attacked LLMs can cause persuasive generation “Here is a link…:phishing.website…” without realizing the phishing risks to the user. General backdoor or poisoning attacks teach a LLM to perform differently for triggered inputs in general tasks like sentiment classification. In contrast, our Trojan exploits ignorance of a LLM to human value judgement on a Trojan knowledge injection.
Potential Social Impacts. Our study has meaningful takeaways for policymakers, professionals and the general public. First, professional LLM users and developers should be vigilant when using a third-party adapter. A good practice for reducing the risk of malware is to only install applications from trusted sources. Similarly, it is important to avoid using adapters or models shared by unknown developers. Meanwhile, LLM agents should be placed under sufficient access control, by either executing in a sandbox or by scrutinizing the LLM through an output filter.
Second, LLMs, IT datasets and adapters should be accompanied with identities [84] to ensure traceability. In this way, once some vulnerability is found, the community will be notified and affected models can be removed from deployment. This strategy is akin to the suggestion we received from Hugging Face to report vulnerabilities on the bug bounty platform “huntr”. The model platform can also provide warrants to ensure the security of shared models by, for example, proof-of-learning [85]. Another solution is establishing effective licensing and a governance system for generative models [86].
Third, for non-professional users, our study reiterates the importance of being careful about information from, not only the LLM, as LLMs can exactly follow an adversary order under certain conditions, but also the Internet. To counter misinformation, one should make critical decisions with information from multiple sources. Additionally, it is important to promptly report encountered misbehavior to the LLM administrators, which allows the malfunctioned model (either attacked or not) to be fixed swiftly.
Limitations. Our attacks obtain a lower attack success rate under low injection rates. This hinders attack effectiveness if the adversary can only poison training data with polished. For example, a malicious doctor, with the improved attack, can modify his or her prescriptions (occupying a small proportion in the training data) to recommend drugs of interest. Further, our attacks rely on a fixed trigger for activation regardless of the input context (e.g., chat history). This increases detectability when the system is operational: the operator can unload adapters after receiving a bug report about abnormal outputs from vigilant users. One possible direction is designing a context-aware Trojan to increase the credibility of misinformation.
Future Work. To defend the Trojan threat in adapters, we plan to develop an efficient evolution for vulnerable trigger scanning. Although this approach is similar to LLM red-teaming, existing red-teamed LLMs are still vulnerable to our attack. We found our baseline attack can still be highly successful on the red-teamed LLaMA-2 and LLaMA-3 models. The concurrent work [74] also confirms that conventional mitigations, such as safety alignment, cannot remove a deceptive backdoor in a LLM. This is likely because the Trojan target is not considered during red-teaming.
Our exposed threat should be effective for foundation models of other modalities. For example, text-to-image models like Stable Diffusion (SD) heavily rely on LoRAs for model personalization. A compromised adapter of SD can produce harmful content for sensitive topics, causing significant consequences once being deployed online. Hence, we plan to extend our work to other foundation models.
Acknowledgment
The authors from Shanghai Jiao Tong University were partially supported by the National Natural Science Foundation of China (No. 62325207, 62132013, 62302298). Minhui Xue is supported in part by Australian Research Council (ARC) DP240103068 and in part by CSIRO – National Science Foundation (US) AI Research Collaboration Program. We would like to thank anonymous reviews for their insightful feedback. We also thank Yanzhu Guo and Yiming Wang for their discussion at early stage of the project. Haojin Zhu (zhu-hj@sjtu.edu.cn) is the corresponding author